An object is thrown directly downward from the top of a very tall building. The speed of the object just as it is released is 29.3 m/s. After being thrown, the object falls freely due to gravity. Neglect air resistance and calculate the distance, in meters which the object covers between times t1 = 1.77 s and t2 = 5.02 s after it is thrown.
Answer #1

a = 9.4 m/s 2 = 29-3 x 1-17 +1 x g,ex (1-77) - S1 86 Is 1s S7.2 m t2=$-02 sec S2= 270.56M distence Caves Letvcen ts 2 270-SC-